# PAD: Personalized Alignment of LLMs at Decoding-Time

### v5

@@ Abstract
Aligning with personalized preferences, which vary significantly across cultural, educational, and political differences, poses a significant challenge due to the computational costs and data demands of traditional alignment methods. In response, this paper presents Personalized Alignment at Decoding-time (PAD), a novel framework designed to align LLM outputs with diverse personalized preferences during the inference phase, eliminating the need for additional training. By introducing a unique personalized reward modeling strategy, this framework decouples the text generation process from personalized preferences, facilitating the generation of generalizable token-level personalized rewards. The PAD algorithm leverages these rewards to guide the decoding process, dynamically tailoring the base model’s predictions to personalized preferences. Extensive experimental results demonstrate that PAD not only outperforms existing training-based alignment methods in terms of aligning with diverse preferences but also shows significant generalizability to preferences unseen during training and scalability across different base models. This work advances the capability of LLMs to meet user needs in real-time applications, presenting a substantial step forward in personalized LLM alignment.
@@ 1 Introduction
Recent advancements have demonstrated success in aligning language models with human preferences and values (Stiennon et al., 2020; Bai et al., 2022; Ouyang et al., 2022; Achiam et al., 2023). Representative methods such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) and DPO (Rafailov et al., 2024b) typically optimize a policy model with training signals from an explicit or implicit reward model. The reward model captures the ‘general’ human preferences or values from human feedback. However, in this pluralistic world, users’ preferences can diverge significantly based on their different cultures, educational backgrounds, religions, and political stands (Gordon et al., 2022; Sorensen et al., 2024b; Jang et al., 2023; Cheng et al., 2023). Furthermore, even for the same person, the preference of a particular LLM response can vary when the application scenario changes. Hence, there always exists a proportion of human preferences that cannot be unified by the general preference, also known as personalized preferences, which current alignment frameworks struggle to align with due to the need for high-quality datasets and substantial computational costs in policy optimization.
How can we align with personalized preferences without the need for additional data collection and policy training? In this paper, we introduce Personalized Alignment at Decoding-time (PAD), which aims to align LLM outputs with diverse personalized preferences during the inference phase without requiring additional training. To achieve this, we first propose a personalized reward modeling strategy, which decouples the text generation process (modeled as a Markov Decision Process) from personalized preferences, thereby enabling the acquisition of generalizable token-level personalized rewards. Based on this, we then formulate a personalized reward model (PRM). During decoding, the PRM scores the base model’s top-K predictions at each token generation step based on the current generation and personalized preferences. Finally, this score is combined with standard decoding likelihoods to adjust the base model’s predictions. The advantages of PAD are as follows: (1) It requires only a single policy model (i.e., the base model) aligned with general preferences (General Policy), eliminating the need for training additional policy models (Training-free). (2) It utilizes only a single reward model (Single Reward). (3) It does not require pre-defined personalized preferences to generalize to preferences not seen during the training phase (Generalizability). A checklist of PAD’s advantages over previous methods is presented in Table 1.
@@
  • We propose a novel personalized reward modeling strategy that decouples the dynamics of text generation from personalized preferences. This strategy enables the acquisition of generalizable token-level personalized rewards with a single personalized reward model (PRM).
@@ 2 Related Works
Large language model alignment aims to align LLMs with human preferences. The mainstream methods are training-based. A common approach involves using an RLHF (Reinforcement Learning with Human Feedback) framework Christiano et al. (2017); Bai et al. (2022) where a reward model is trained based on human feedback, and Proximal Policy Optimization (PPO) (Schulman et al., 2017) is employed to derive the aligned policy model. Recent efforts explore alternatives to enhance stability and reduce resources. Notably, DPO (Rafailov et al., 2024b) leverages the Bradley-Terry assumption (Bradley & Terry, 1952) for direct optimization of the preference-based objective. Decoding-time alignment offers an alignment paradigm that does not require expensive RL training (Han et al., 2024; Liu et al., 2024). Controlled Decoding (CD) (Mudgal et al., 2023) utilizes a prefix scorer module trained to assess value functions for rewards, allowing controlled generation from a frozen base model. ARGS (Khanov et al., 2024) proposed using a reward signal to adjust probabilistic predictions, thereby generating semantically aligned texts. DeAL (Huang et al., 2024) focusing on heuristic-guided searches to better meet diverse alignment objectives.
@@
As humans exhibit diverse preferences and values for a single task, it is essential to align large language models (LLMs) to users’ personalized preferences (Kirk et al., 2023; Sorensen et al., 2023; 2024c; Yao et al., 2023; Kirk et al., 2024; Zhong et al., 2024; Han et al., 2024). One line of work achieves joint optimization for different personalized preferences by defining a reward function with multiple dimensions and performing policy optimization (Zhou et al., 2023; Wang et al., 2024a; b; Guo et al., 2024; Yang et al., 2024b; Chakraborty et al., 2024; Sun et al., 2024; Li et al., 2024). Additionally, some approaches involve merging model parameters or predictions trained for each dimension to accommodate the diverse combinations expressed by those dimensions (Jang et al., 2023; Rame et al., 2024; Park et al., 2024; Shi et al., 2024). Lastly, prompt-based methods align personalized user preferences by designing diverse prompts or post-processing techniques (Yang et al., 2024a; Lee et al., 2024; Hwang et al., 2023b; Jafari et al., 2024). The work most similar to ours is MOD (Shi et al., 2024), which achieves flexible trade-offs and optimization across multiple objectives by linearly combining predictions from different policy models at decoding time to output the next token. However, MOD still requires training base models for different preferences, making it difficult to scale to a large number of personalized preferences.
@@ 3.2 Personalized alignment at Decoding-time
$\mathcal{L}_{\text{PRM}}(\pi_{\theta},D)=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim D}\left[\log\sigma\left(\mathbf{w}_{p}^{\top}(\sum_{t=1}^{T}\beta\log\frac{\hat{\pi}_{\theta}(\mathbf{a}^{w}_{t}|\mathbf{s}^{w}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}^{w}_{t}|\mathbf{s}^{w}_{t})}-\sum_{t=1}^{T}\beta\log\frac{\hat{\pi}_{\theta}(\mathbf{a}^{l}_{t}|\mathbf{s}^{l}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}^{l}_{t}|\mathbf{s}^{l}_{t})})\right)\right].$ (9)
@@
The detailed derivations are provided in the appendix C.1. It can be observed that the learned reward function can serve as an optimal advantage function. Note that unlike RLHF framework that directly models the language model as policy, our PAD policy $\pi_{\textsc{PAD}}$ is different from the base language model $\pi_{LM}$, as well as the personalized reward model $\pi_{\theta}$.
Figure 1: An illustration of the inference phase for personalized alignment at decoding-time (PAD) with optimized personalized reward model (PRM). Given the personalized preference and the current context, we first calculate the probability distribution of the base model for the next token. Then, we calculate the reward from PRM combining features of current state and personalized weight. Finally, the next token can be selected based on the weighted scores.
@@ 3.3 Generalized personalized alignment
$Q_{n+1}^{\pi^{*}}(p_{n+1},\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(p,\mathbf{s},\mathbf{a})\leq|H|\left(\phi_{\max}\min_{j}\|\mathbf{w}_{{n+1}}-\mathbf{w}_{j}\|\right).$ (12)
This term is a multiple of the distance between $\mathbf{w}_{i}$ and the closest $\mathbf{w}_{j}$ for which we have already computed a policy. The formula formalizes the intuition that if the agent has previously solved a similar task, it should perform well on task $\mathbf{w}_{i}$. The proofs of Theorem 1 are in the Appendix C.2.
@@ 3.4 Practical Implementations
In this section, we introduce the practical implementation of our Personalized Alignment at Decoding-time (PAD), which includes the optimization of the personalized reward model (PRM) and the inference-time guided decoding with token-level personalized reward.
@@
As previously mentioned, we can decouple personalized preferences from the MDP process during the personalized alignment procedure. Thus, we consider utilizing an LLM-based model as a personalized reward model (PRM) $\pi_{\theta}$ to independently predict the features $\phi(\mathbf{s},\mathbf{a})$ and preferences $\mathbf{w}_{p}$. For simplicity, we employ a single backbone to predict the embeddings for both $\phi(\mathbf{s},\mathbf{a})$ and $p$, and then utilize an additional value head to predict $\mathbf{w}_{p}$. Optimization occurs in two stages with Equation 9. In the first stage, we fix $\mathbf{w}_{p}$ as a unit vector and optimize the backbone to learn general features. In the second stage, we freeze the backbone and only optimize the value head for $\mathbf{w}_{p}$ to learn different user preferences. The optimized PRM is denoted as $\pi^{*}_{\theta}$.
@@
The inference phase of PAD is shown in Figure 1. Given the personalized preference $p$ and the current context $s=(x,y_{<t})$ at step $t$, we first calculate the predicted probability distribution of the base model $\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})$ for the next token $\mathbf{a}$. Then we calculate the reward from PRM by $R(p,\mathbf{s}_{t},\mathbf{a})=\mathbf{w}_{p}^{\top}\phi(\mathbf{s}_{t},\mathbf{a})=\mathbf{w}_{p}^{\top}\log(\hat{\pi}^{*}_{\theta}(\mathbf{a}|\mathbf{s}_{t})/{\hat{\pi}_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})})$ for the token. Finally, $\mathbf{a}_{t}$ can be selected based on their weighted scores:
@@
Eq. 13 is equivalent to Eq. 11, with proofs in the Appendix C.1. Note that $\beta$ in Eq. 13 is also treated as a weight hyperparameter for simplicity, which slightly differs from its previous definition.
@@ 4.1.1 PAD Setup
During the development of our personalized reward model, we utilized datasets from multiple sources including Ultrafeedback (Cui et al., 2023), HelpSteer2 (Wang et al., 2024c), Rewards-in-Context (Yang et al., 2024b), and SafeRLHF (Dai et al., 2023). To simulate personalized preferences, synthetic user prompts are prepended to instructions, following Jang et al. (2023). We employ the Llama-3-8B model (AI@Meta, 2024) as our backbone, and append a linear layer directly following the embeddings, featuring an output dimension of 4096. Comprehensive details on the datasets, the pre-processing process and the implementations are documented in Appendix A.1. During the decoding phase, we utilize greedy decoding with top-$k$ candidates. We restrict the maximum lengths of the initial prompt and subsequent generations to 2,048 and 128 tokens, respectively. The hyperparameters, specifically $\beta=1.0$ and $k=10$, are optimized to maximize the average reward performance observed in our validation datasets. An exhaustive analysis of the decoding strategies and hyperparameter settings is provided in Section 4.3.
@@ 4.1.2 Evaluation Setup
We utilize two evaluation datasets. The P-Soups (Jang et al., 2023) evaluation dataset has been filtered and modified based on the Koala evaluation by Jang et al. (2023). The HelpSteer2 (Wang et al., 2024c) (validation split) dataset is a multi-aspect alignment dataset comprising 1,000 prompts. In our evaluation setup, we initially focus on alignment the three pre-defined dimensions: “harmless”, “helpful”, and “humor”, following previous works (Yang et al., 2024b; a; Shi et al., 2024). Additionally, to assess the scalability of the Personalized Alignment Dataset (PAD), we simulate users each having unique personalized preferences drawn from several preference dimensions, collectively forming 12 preference combinations. For personalized alignment, we primarily utilize LLama-3-8B-SFT (AI@Meta, 2024; Meng et al., 2024) as the base language model. Additional experiments are conducted on Gemma (Team, 2024), Mistral-7B-SFT (Jiang et al., 2023; Tunstall et al., 2023), and Llama-2 (Touvron et al., 2023) to test scalability.
@@
For the evaluation of “harmless”, “helpful”, and “humor” dimensions, we integrate three open-source reward models available on Huggingface, as delineated by Yang et al. (2024b; a). Additionally, we employ the ArmoRM (Wang et al., 2024b), a multi-dimension reward model known for its state-of-the-art performance on the Reward-Bench benchmark (Lambert et al., 2024). For these reward models, we report their scores assessing LLM responses from various perspectives. Furthermore, our evaluation leverages GPT-4, a widely recognized tool in previous studies (Khanov et al., 2024; Yang et al., 2024a), to conduct the judgments towards certain personalized preference and report the win rate. Comprehensive details on the evaluation metrics, reward models and GPT-4 judgments are provided in Appendix A.3.
@@
We compare PAD with 9 personalized alignment or multi-objective alignment methods as delineated in Table 1. MORLHF (Li et al., 2020) optimizes for the weighted multi-objective reward function using PPO. MODPO (Zhou et al., 2023) integrates language modeling with reward modeling, training models to combine all objectives with specific weights. Personalized soups (Jang et al., 2023) first optimizes multiple policy models and then merges the parameters of the policy models according to preference. Rewarded soup (Rame et al., 2024) first trains multiple specializing networks independently, and then interpolates their weights linearly. Rewards-in-Context (RiC) (Yang et al., 2024b) conditions foundation model responses on multiple rewards in its prompt and uses supervised fine-tuning for alignment. MetaAligner (Yang et al., 2024a) trains an additional model to perform conditional weak-to-strong correction. Preference prompting (Jang et al., 2023) simply prompts for preferences without any additional training. MOD (Shi et al., 2024) first trains multiple specializing networks and performs linear combination of their predictions. Args (Khanov et al., 2024) is a reward-guided decoding-time alignment framework. As Args is originally designed for general preference, we use ArmoRM (Wang et al., 2024b) to provide rewards.

### v7

@@ Abstract
Aligning with personalized preferences, which vary significantly across cultural, educational, and political differences, poses a significant challenge due to the computational costs and data demands of traditional alignment methods. In response, this paper presents Personalized Alignment at Decoding-time (PAD), a novel framework designed to align LLM outputs with diverse personalized preferences during the inference phase, eliminating the need for additional training. By introducing a unique personalized reward modeling strategy, this framework decouples the text generation process from personalized preferences, facilitating the generation of generalizable token-level personalized rewards. The PAD algorithm leverages these rewards to guide the decoding process, dynamically tailoring the base model’s predictions to personalized preferences. Extensive experimental results demonstrate that PAD not only outperforms existing training-based alignment methods in terms of aligning with diverse preferences but also shows significant generalizability to preferences unseen during training and scalability across different base models. This work advances the capability of LLMs to meet user needs in real-time applications, presenting a substantial step forward in personalized LLM alignment. Our model and code are available here.
@@ 1 Introduction
Recent advancements have demonstrated success in aligning language models with human preferences and values (Stiennon et al., 2020; Bai et al., 2022; Ouyang et al., 2022; Achiam et al., 2023). Representative methods such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) and DPO (Rafailov et al., 2024b) typically optimize a policy model with training signals from an explicit or implicit reward model. The reward model captures the ‘general’ human preferences or values from human feedback. However, in this pluralistic world, users’ preferences can diverge significantly based on their different cultures, educational backgrounds, religions, and political stands (Gordon et al., 2022; Sorensen et al., 2024a; Jang et al., 2023; Cheng et al., 2023). Furthermore, even for the same person, the preference of a particular LLM response can vary when the application scenario changes. Hence, there always exists a proportion of human preferences that cannot be unified by the general preference, also known as personalized preferences, which current alignment frameworks struggle to align with due to the need for high-quality datasets and substantial computational costs in policy optimization.
How can we align with personalized preferences without the need for additional data collection and policy training? In this paper, we introduce Personalized Alignment at Decoding-time (PAD), which aims to align LLM outputs with diverse personalized preferences during the inference phase without requiring additional training. To achieve this, we first propose a personalized reward modeling strategy, which decouples the text generation process (modeled as a Markov Decision Process) from personalized preferences, thereby enabling the acquisition of generalizable token-level personalized rewards. Based on this, we then formulate a personalized reward model (PersRM). During decoding, the PersRM scores the base model’s top-K predictions at each token generation step based on the current generation and personalized preferences. Finally, this score is combined with standard decoding likelihoods to adjust the base model’s predictions. The advantages of PAD are as follows: (1) It requires only a single policy model (i.e., the base model) aligned with general preferences (General Policy), eliminating the need for training additional policy models (Training-free). (2) It utilizes only a single reward model (Single Reward). (3) It does not require pre-defined personalized preferences to generalize to preferences not seen during the training phase (Generalizability). A checklist of PAD’s advantages over previous methods is presented in Table 1.
@@
  • We propose a novel personalized reward modeling strategy that decouples the dynamics of text generation from personalized preferences. This strategy enables the acquisition of generalizable token-level personalized rewards with a single personalized reward model (PersRM).
@@ 2 Related Works
Large language model alignment aims to align LLMs with human preferences. A common approach involves using an RLHF (Reinforcement Learning with Human Feedback) framework Christiano et al. (2017); Bai et al. (2022) where a reward model is trained based on human feedback, and Proximal Policy Optimization (PPO) (Schulman et al., 2017) is employed to derive the aligned policy model. Notably, decoding-time alignment offers an alignment paradigm that does not require expensive RL training (Mudgal et al., 2023; Khanov et al., 2024; Han et al., 2024; Liu et al., 2024; Huang et al., 2024). Controlled Decoding (CD) (Mudgal et al., 2023) utilizes a prefix scorer module trained to assess value functions for rewards, allowing controlled generation from a frozen base model. ARGS (Khanov et al., 2024) proposed using a reward signal to adjust probabilistic predictions, thereby generating semantically aligned texts. DeAL (Huang et al., 2024) focusing on heuristic-guided searches to better meet diverse alignment objectives.
@@
As users exhibit diverse preferences and values for a single task, it is essential to align LLMs to personalized preferences (Kirk et al., 2023; Sorensen et al., 2023; 2024b; Yao et al., 2023; Kirk et al., 2024; Zhong et al., 2024; Han et al., 2024). One line of work achieves joint optimization for different personalized preferences by defining a reward function with multiple dimensions and performing policy optimization (Zhou et al., 2023; Wang et al., 2024a; b; Guo et al., 2024; Yang et al., 2024b; Chakraborty et al., 2024; Sun et al., 2024; Li et al., 2024). Additionally, some approaches merge model parameters or predictions for each dimension to handle their diverse combinations (Jang et al., 2023; Rame et al., 2024; Park et al., 2024; Shi et al., 2024). Lastly, prompt-based methods align personalized preferences by designing diverse prompts or post-processing techniques (Yang et al., 2024a; Lee et al., 2024; Hwang et al., 2023; Jafari et al., 2024).
@@ 3.2 Personalized alignment at Decoding-time
$\mathcal{L}_{\text{PersRM}}(\pi_{\theta},D)=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim D}\left[\log\sigma\left(\mathbf{w}_{p}^{\top}(\sum_{t=1}^{T}\beta\log\frac{\hat{\pi}_{\theta}(\mathbf{a}^{w}_{t}|\mathbf{s}^{w}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}^{w}_{t}|\mathbf{s}^{w}_{t})}-\sum_{t=1}^{T}\beta\log\frac{\hat{\pi}_{\theta}(\mathbf{a}^{l}_{t}|\mathbf{s}^{l}_{t})}{\hat{\pi}_{\text{ref}}(\mathbf{a}^{l}_{t}|\mathbf{s}^{l}_{t})})\right)\right].$ (9)
@@
The detailed derivations are provided in the appendix D.1. It can be observed that the learned reward function can serve as an optimal advantage function. Note that unlike RLHF framework that directly models the language model as policy, our PAD policy $\pi_{\textsc{PAD}}$ is different from the base language model $\pi_{LM}$, as well as the personalized reward model $\pi_{\theta}$.
Figure 1: An illustration of the inference phase for personalized alignment at decoding-time (PAD) with optimized personalized reward model (PersRM). Given the personalized preference and the current context, we first calculate the probability distribution of the base model for the next token. Then, we calculate the reward from PersRM combining features of current state and personalized weight. Finally, the next token can be selected based on the weighted scores.
@@ 3.3 Generalized personalized alignment
$Q_{n+1}^{*}(p_{n+1},\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(p,\mathbf{s},\mathbf{a})\leq|H|\left(\phi_{\max}\min_{j}\|\mathbf{w}_{{n+1}}-\mathbf{w}_{j}\|\right).$ (12)
This term is a multiple of the distance between $\mathbf{w}_{{n+1}}$ and the closest $\mathbf{w}_{j}$ for which we have already computed a policy. The formula formalizes the intuition that if PAD has previously aligned to a similar personalized preference $\mathbf{w}_{j}$, it should align well on the new preference $\mathbf{w}_{{n+1}}$ Barreto et al. (2017). The proofs of Theorem 1 are in the Appendix D.2.
@@ 3.4 Practical Implementations
In this section, we introduce the practical implementation of our Personalized Alignment at Decoding-time (PAD), which includes the optimization of the personalized reward model (PersRM) and the inference-time guided decoding with token-level personalized reward.
@@
As previously mentioned, we can decouple personalized preferences from the MDP process during the personalized alignment procedure. Thus, we consider utilizing an LLM-based model as a personalized reward model (PersRM) $\pi_{\theta}$ to independently predict the features $\phi(\mathbf{s},\mathbf{a})$ and preferences $\mathbf{w}_{p}$. For simplicity, we employ a single backbone to predict the embeddings for both $\phi(\mathbf{s},\mathbf{a})$ and $p$. $\pi_{\theta}$ is initialized using the backbone, also referred to as the reference model $\pi_{\text{ref}}$. Subsequently, it employs an additional value head to predict $\mathbf{w}_{p}$. Optimization occurs in two stages with Equation 9. In the first stage, we fix $\mathbf{w}_{p}$ as a unit vector and optimize the backbone to learn general features. In the second stage, we freeze the backbone and only optimize the value head for $\mathbf{w}_{p}$ to learn different user preferences. The optimized PersRM is denoted as $\pi^{*}_{\theta}$.
@@
The inference phase of PAD is shown in Figure 1. Given the personalized preference $p$ and the current context $\mathbf{s}_{t}=(\mathbf{x},\mathbf{y}_{<t})$ at step $t$, we first calculate the predicted probability distribution of the base model $\pi_{LM}(\mathbf{a}|\mathbf{s}_{t})$ for each next token candidate $\mathbf{a}$. Then we calculate the reward from PersRM by $R(p,\mathbf{s}_{t},\mathbf{a})=\mathbf{w}_{p}^{\top}\phi(\mathbf{s}_{t},\mathbf{a})=\mathbf{w}_{p}^{\top}\log(\hat{\pi}^{*}_{\theta}(\mathbf{a}|\mathbf{s}_{t})/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\hat{\pi}_{\text{ref}}(\mathbf{a}|\mathbf{s}_{t})}})$ for these tokens. Finally, the next token $\mathbf{a}_{t}$ can be selected based on their weighted scores:
@@
Eq. 13 is equivalent to Eq. 11, with proofs in the Appendix D.1. Note that $\beta$ in Eq. 13 is also treated as a weight hyperparameter for simplicity, which slightly differs from its previous definition.
@@ 4.1.1 PAD Setup
During the development of our personalized reward model, we utilized datasets from multiple sources including HelpSteer2 (Wang et al., 2024c), Rewards-in-Context (Yang et al., 2024b), and SafeRLHF (Dai et al., 2023). To simulate personalized preferences, synthetic user prompts are prepended to instructions, following Jang et al. (2023). We employ the Llama-3-8B model (AI@Meta, 2024) as our backbone, and append a linear layer directly following the embeddings, featuring an output dimension of 4096. Comprehensive details on the datasets, the pre-processing process and the implementations are documented in Appendix A. During the decoding phase, we utilize greedy decoding with top-$k$ candidates. We restrict the maximum lengths of the initial prompt and subsequent generations to 2,048 and 128 tokens, respectively. The hyperparameters, specifically $\beta=1.0$ and $k=10$, are optimized to maximize the average reward performance observed in our validation datasets. Analysis of the decoding strategies and hyperparameter settings is provided in Section 4.3.
@@ 4.1.2 Evaluation Setup
We utilize two evaluation datasets. The P-Soups (Jang et al., 2023) evaluation dataset has been filtered and modified based on the Koala evaluation by Jang et al. (2023). The HelpSteer2 (Wang et al., 2024c) (validation split) dataset is a multi-aspect alignment dataset comprising 1,000 prompts. In our evaluation setup, we initially focus on alignment the three pre-defined dimensions: ‘harmless’, ‘helpful’, and ‘humor’, following previous works (Yang et al., 2024b; a; Shi et al., 2024). Additionally, to assess the scalability of the Personalized Alignment Dataset (PAD), we simulate users each having unique personalized preferences drawn from several preference dimensions, collectively forming 12 preference combinations. For personalized alignment, we primarily utilize Llama-3-8B-SFT (AI@Meta, 2024; Meng et al., 2024) as the base language model. Additional experiments are conducted on Gemma (Team, 2024), Mistral-7B-SFT (Jiang et al., 2023; Tunstall et al., 2023), and Llama-2 (Touvron et al., 2023) to test scalability.
@@
For the evaluation of ‘harmless’, ‘helpful’, and ‘humor’ dimensions, we integrate three open-source reward models available on Huggingface, as delineated by Yang et al. (2024b; a). Additionally, we employ the ArmoRM (Wang et al., 2024b), a multi-dimension reward model known for its state-of-the-art performance on the Reward-Bench benchmark (Lambert et al., 2024). For these reward models, we report their scores assessing LLM responses from various perspectives. Furthermore, our evaluation leverages GPT-4, a widely recognized tool in previous studies (Khanov et al., 2024; Yang et al., 2024a), to conduct the judgments towards certain personalized preference and report the win rate. Comprehensive details on the evaluation metrics, reward models and GPT-4 judgments are provided in Appendix B.2.
@@
We compare PAD with 9 personalized alignment or multi-objective alignment methods as delineated in Table 1. MORLHF (Li et al., 2020) optimizes for the weighted multi-objective reward function using PPO. MODPO (Zhou et al., 2023) integrates language modeling with reward modeling, training models to combine all objectives with specific weights. Personalized soups (Jang et al., 2023) first optimizes multiple policy models and then merges the parameters of the policy models according to preference. Rewarded soup (Rame et al., 2024) first trains multiple specializing networks independently, and then interpolates their weights linearly. Rewards-in-Context (RiC) (Yang et al., 2024b) conditions foundation model responses on multiple rewards in its prompt and uses supervised fine-tuning for alignment. MetaAligner (Yang et al., 2024a) and Aligner (Ji et al., 2024) train an additional model to perform conditional weak-to-strong correction. Preference prompting (Jang et al., 2023) simply prompts for preferences without any additional training. MOD (Shi et al., 2024) first trains multiple specializing networks and performs linear combination of their predictions.